1. The fourth term of a geometric sequence is 2 and the seventh term is 16; determine:
(a) the first two terms
(b) the sum of the first 20 terms (two decimals)

Answer by passionn(5)

Que 1a) Determining first two terms
a, ar^2, ar^3,.........ar^(n-1)
ar^3=2
ar^6=16
therefore 1st term is calculated as follows:
ar^6/ar^3= 16*1/2= r^3= 8
since ar^3=2
a(8)/8=2/8=a=1/4
1st term is 0.25
Second term is:
ar^(n-1)= 0.25(2)^(2-1)= 0.5
Que 1b) Determining the sum of the first 20 terms (two decimals)
Sn= a(r^n-1)/(r-1)
S20= 0.25(2^20-1)/(2-1)
S20= 262143.75